# Multivariate Diffusion Transformer with Decoupled Attention for

### v1

@@ 1 Introduction
High-fidelity facial synthesis (ning2023multi; chen2024dreamidentity; melnik2024face; wang2025facea) holds transformative potential across digital identity systems, adversarial security, and immersive media applications (terhorst2023qmagface; rehaan2024face; diao2025ft2tf). While recent advances in generative modeling, such as GANs and diffusion models, have enabled photorealistic facial generation, deploying these technologies in practice remains hindered by in fine-grained controllability, posing challenges in precisely generating facial images that align with user expectations.
Modern conditional generation frameworks have transitioned from unimodal to multimodal paradigms to tackle these challenges. Initially, text-driven methods constructed end-to-end pipelines to transform semantic descriptions into facial attributes (kim2022diffusionclip; zhu_grainedclip_2023). However, they encountered difficulties in achieving precise spatial localization (e.g., pose, geometry) due to the inherent ambiguity of language. Subsequently, visual modalities such as semantic masks (chen2022sofgan; ergasti2024controllable) were incorporated to enforce geometric constraints, but this often came at the cost of reduced semantic versatility in attributes like age, gender, or complexion. This dichotomy has spurred research into multimodal collaborative facial synergies (meng2025mm2latent; sowmya2024generative; kim2024diffusion) that leverage the expressive capacity of natural language for high-level attributes and the precision of visual cues for low-level spatial control.
Existing state-of-the-art multimodal facial synthesis methods generally adhere to two primary architectural paradigms: (1) Generative Adversarial Network (GAN)-based frameworks (du2023pixelface+; kim2024diffusion; meng2025mm2latent), such as StyleGAN variants, which project heterogeneous modality conditions into a latent space via multiple encoding networks for subsequent fusion; and (2) diffusion models (nair2023unite; huang2023collaborative; peng2024controlnext; kim2024diffusion), which integrate unimodal expert diffusion models through dynamic weighting during the inference process. However, these approaches typically process multiple modalities in isolation or employ simplistic feature fusion strategies, resulting in suboptimal multimodal feature interaction—particularly in scenarios that demand collaborative optimization between visual masks and textual descriptions.
@@ 2.1 Diffusion Models for Image Generation
Recent advancements in generative modeling have witnessed a paradigm shift from GAN-based architectures (you2022application; du2023pixelface+; che2024assessing) to diffusion models (po2024state; he2025diffusion; chang2025design; huang2025diffusion), driven by their superior training stability and output diversity. Early diffusion models like DDPM (ho2020denoising) and DDIM (song2020denoising) established foundational sampling pipelines, while subsequent works focused on computational efficiency (rombach2022high; xue2024accelerating; chadebec2025flash) and conditional control (podell2023sdxl; yoon2023score; peng2024controlnext). Latent diffusion models (LDM) (rombach2022high) reduced computational costs by operating in compressed latent spaces, enabling high-resolution synthesis. ControlNet (zhang2023adding), ControlNeXt (peng2024controlnext), and T2I-Adapter (mou2024t2i) further enhanced controllability by introducing external conditioning networks, though their multimodal integration is still confined to the addition of independent feature maps.
The integration of vision transformers has driven recent breakthroughs, where diffusion transformers (DiTs) (peebles2023scalable; wang2025lavin) demonstrate superior performance over U-Net architectures (ronneberger2015u) through global attention mechanisms, with representative works including FLUX.1 (flux2024), OminiControl (tan2024ominicontrol; tan2025ominicontrol2), and OmniGen (xiao2025omnigen). However, these methods often incur significantly higher computational costs.
@@ 2.2 Facial Image Synthesis with Multimodal Conditioning
Facial synthesis techniques (kim2023dcface; melnik2024face; song2025attridiffuser) have evolved from unimodal control paradigms to multimodal control paradigms, with recent research efforts focusing on the collaborative integration of various modalities, such as text and semantic masks (du2023pixelface+; meng2025mm2latent). Methods like TediGAN (xia2021tedigan) and MM2Latent (meng2025mm2latent) leverage the disentangled properties of the $\mathcal{W}$ latent space in StyleGANs to achieve multimodal-driven facial synthesis by fusing latent vectors from multiple modalities. In contrast, diffusion-based approaches, such as GCDP (park2023learning), have aimed to enhance semantic alignment in synthesized faces; however, they are limited by the restrictive assumptions of joint Gaussian distributions. Additionally, methods like UaC (nair2023unite) and CoDiffusion (huang2023collaborative) dynamically integrate multiple unimodal expert models during inference, which results in a more than two-fold increase in computational cost. Meanwhile, DDG (kim2024diffusion) combines diffusion models with GANs, utilizing ControlNet (zhang2023adding; peng2024controlnext) for semantic mask and text multimodal feature fusion. Still, the expressiveness of the StyleGAN latent space representation constrains the quality of synthesized faces.
@@ 3.1 Preliminaries
Diffusion transformer (DiT) (peebles2023scalable) successfully breaks through the limitations of local perception by adopting a pure transformer architecture to replace the traditional U-Net backbone network. This achieves an integration of global dependency modeling and multimodal information fusion, thereby significantly enhancing the model’s capability to model semantic consistency in complex scenarios. At time step $t$, the model utilizes stacked tansformer blocks to synchronously process noisy image tokens $\mathbf{X}_{t}\in\mathbb{R}^{N\times d}$ and text condition tokens $\mathbf{C}_{T}\in\mathbb{R}^{L\times d}$. Each transformer block incorporates a multi-head attention computation via Equation 1, ensuring in-depth interaction and fusion between visual and semantic information.
@@
Rotary Positional Embedding. To overcome the translation invariance limitation of traditional absolute position encoding, DiT introduces Rotary Position Encoding (RoPE) (heo2024rotary) to explicitly model spatial relationships. Each visual token at position $(i,j)$ is modulated via:
@@ 3.2 Model Design
Figure 1 illustrates the overall framework of the proposed MDiTFace method, which aims to synthesize high-fidelity facial images under the collaborative constraints of textual descriptions and semantic masks. Built upon the FLUX.1 text-to-image generation model (flux2024; labs2025flux1kontextflowmatching), MDiTFace employs a unified tokenization strategy to process heterogeneous mask-text conditional inputs, projecting them through modality-specific embedders into a shared latent space where the noise image tokens $\mathbf{X}_{t}$ resides, thereby eliminating representation disparities across heterogeneous modalities. To achieve synchronous processing and achieve a thorough fusion of multiple conditional token sequences, which include mask tokens $C_{M}$ and text tokens $C_{T}$, we have conducted a comprehensive redesign and subsequent implementation of the multivariate transformer block. These blocks incorporate an innovatively designed decoupled attention mechanism that enables flexible interaction among multimodal tokens while partitioning internal computations into dynamic and static pathways. By eliminating the implicit dependency between mask tokens $C_{M}$ and temporal embeddings, the relevant features of the static pathway can be cached after the initial computation and reused across denoising steps, thereby significantly reducing redundant computational overhead.
@@
However, experimental results indicate that this simple fusion strategy leads to poor mask-condition consistency in synthesized facial images (see Table 3 for details). To address this issue, our proposed MDiTFace method employs a unified tokenization strategy to process multimodal input conditions: The mask is encoded by reusing the VAE encoder, preserving spatial position information while maintaining architectural simplicity. The text is encoded using the pre-trained T5 encoder (raffel2020exploring) to capture semantic richness. Subsequently, the encoded features from both modalities are projected into a latent space shared with the noisy image tokens $\mathbf{X}_{t}$ via modality-specific embedders, bridging representation gaps between heterogeneous modalities. This transformation is formalized in Equation 4:
@@ 3.4 Training Strategy
Training Objective. We employ a lightweight adaptation architecture, which only requires low-rank adaptation (LoRA) (hu_lora:_2021) fine-tuning on the visual embedder and some linear layers within the multivariate transformer blocks. This approach results in training fewer than 0.1% additional parameters, thereby significantly reducing the computational cost of training. The model is optimized using the flow matching loss, with the specific formula as follows:
@@ 4 Experiments
Comparisons TediGAN (xia2021tedigan) GAN 0.7130 0.4885 1.562 87.89 23.90 2.46 -0.1446
UaC(nair2023unite) Diffusion 0.6027 0.5761 1.982 84.71 25.52 3.41 -0.4001
PixelFace+ (du2023pixelface+) GAN 0.5720 0.5640 1.273 93.82 26.16 2.61 0.4608
CoDiffusion (huang2023collaborative) Diffusion 0.7381 0.5845 0.734 83.75 24.51 3.22 -0.1265
ControlNeXt (peng2024controlnext) Diffusion 0.8037 0.5469 1.042 91.65 25.88 2.66 0.3675
MM2Latent (meng2025mm2latent) GAN 0.8252 0.5124 1.178 85.05 24.61 3.05 0.2754
@@ 4.1 Experimental Setup
Datasets. MM-CelebA (lee2020maskgan) is a large-scale, multimodal facial synthesis benchmark dataset widely used for evaluation. It comprises 30,000 highly diverse, high-resolution real-world facial images. Each image is accompanied by 10 distinct textual descriptions and a manually annotated semantic mask. The semantic masks cover 19 semantic categories, including major facial components such as hair, skin, eyes, and nose, as well as accessory categories like glasses and clothing. For model training and testing, the MM-CelebA dataset is split into training and test sets in a 9:1 ratio. In addition, to further evaluate the generalization capability of our model, we extended the FFHQ-Text dataset (zhou2021generative) into a mask-text multimodal version, denoted as MM-FFHQ, where the semantic masks were obtained using a pre-trained facial parser (zheng2022general).
Metrics. We employ a range of widely recognized metrics to evaluate the performance of our model, spanning image quality, conditional consistency, and human preference. For image quality,we employ the no-reference TOPIQ metric (chen2024topiq) to evaluate the overall quality of synthesized facial images. Additionally, we utilize the CLIP Maximum Mean Discrepancy (CMMD) to quantify the distributional divergence between synthesized and real facial images in the CLIP feature space (jayasumana2024rethinking), and the Learned Perceptual Image Patch Similarity (LPIPS) to assess perceptual distortion between synthesized and original images (zhang2018unreasonable). For conditional consistency, we measure semantic alignment between text and images using the CLIP score (radford2021learning), evaluate pixel-level alignment between segmentation masks and images via Mask Accuracy, and quantify structural similarity between synthesized and real facial images using the DINO Structure Distance (DSD) (tumanyan2022splicing). Additionally, we employ a pre-trained aesthetic scoring model (xu_imagereward:_2023) to compute the Image Reward Score (IRS), which serves as an indirect proxy for the human-centric visual appeal of the generated images.
Implementation Details. All experiments were conducted on a server equipped with 8 NVIDIA A100 GPUs, each with 80 GB of memory. We employed parameter-efficient LoRA fine-tuning with a rank of 8, and selected Prodigy (mishchenko2024prodigy) as the optimizer. The base learning rate was set to 1.0, and the weight decay coefficient was configured as 0.01. During training, we used a batch size of 16 and optimized the model for 5,000 steps. For inference, we employed the Flow-Matching Euler Discrete scheduler, with a random seed of 42 and 28 sampling steps.
@@ 4.2 Quantitative Evaluation
Comparisons TediGAN (xia2021tedigan) GAN 0.6992 0.6420 1.091 86.81 25.03 3.32 -0.4847
UaC (nair2023unite) Diffusion 0.6106 0.6752 1.796 86.17 26.97 6.38 -0.8851
PixelFace+ (du2023pixelface+) GAN 0.5609 0.6705 1.917 95.02 26.60 3.08 -0.2616
CoDiffusion (huang2023collaborative) Diffusion 0.5035 0.6767 2.178 81.78 23.03 3.85 -1.2101
ControlNeXt (peng2024controlnext) Diffusion 0.7594 0.6165 1.073 86.52 27.80 3.84 -0.4063
MM2Latent (meng2025mm2latent) GAN 0.8198 0.6371 0.437 84.24 26.37 3.52 -0.1197
@@ 4.3 Qualitative Comparison
Figure 3 displays the facial images synthesized by the proposed MDiTFace and competing methods under the same multimodal conditions. It is evident that the facial images synthesized by MDiTFace exhibit higher fidelity. Meanwhile, in terms of fine-grained attributes such as earrings, hair color, and hat shape, MDiTFace can better align with the input conditions, as shown in rows 1-3. This is consistent with the results of quantitative evaluation experiments. In contrast, existing methods suffer from various deficiencies. For instance, methods like UaC (nair2023unite), PixelFace+ (du2023pixelface+), and CoDiffusion (huang2023collaborative) generate facial images with noticeably inadequate fidelity, resulting in unsatisfactory image quality. Other methods, such as TediGAN (xia2021tedigan), ControlNeXt (peng2024controlnext), and MM2Latent (meng2025mm2latent), encounter challenging issues of attribute loss or drift, leading to discrepancies between the synthesized facial images and the expected input conditions. Additionally, when handling the task of facial synthesis under extreme poses, as depicted in row 4, MDiTFace still maintains significant robustness compared to existing methods. More examples can be found in the Supplementary Material.

### v2

@@ 1 Introduction
High-fidelity facial synthesis (Ning et al. 2023; Chen et al. 2024b; Melnik et al. 2024; Wang et al. 2025a) holds transformative potential across digital identity systems, adversarial security, and immersive media applications (Terhörst et al. 2023; Rehaan et al. 2024; Diao et al. 2025). While recent advances in generative modeling, such as GANs and diffusion models, have enabled photorealistic facial generation, deploying these technologies in practice remains hindered by in fine-grained controllability, posing challenges in precisely generating facial images that align with user expectations.
Modern conditional generation frameworks have transitioned from unimodal to multimodal paradigms to tackle these challenges. Initially, text-driven methods constructed end-to-end pipelines to transform semantic descriptions into facial attributes (Kim et al. 2022; Zhu and Mu 2023). However, they encountered difficulties in achieving precise spatial localization (e.g., pose, geometry) due to the inherent ambiguity of language. Subsequently, visual modalities such as semantic masks (Chen et al. 2022; Ergasti et al. 2024) were incorporated to enforce geometric constraints, but this often came at the cost of reduced semantic versatility in attributes like age, gender, or complexion. This dichotomy has spurred research into multimodal collaborative facial synergies (Meng et al. 2025; Sowmya and Meeradevi 2024; Kim et al. 2024) that leverage the expressive capacity of natural language for high-level attributes and the precision of visual cues for low-level spatial control.
Existing state-of-the-art multimodal facial synthesis methods generally adhere to two primary architectural paradigms: (1) Generative Adversarial Network (GAN)-based frameworks (Du et al. 2023; Kim et al. 2024; Meng et al. 2025), such as StyleGAN variants, which project heterogeneous modality conditions into a latent space via multiple encoding networks for subsequent fusion; and (2) diffusion models (Nair et al. 2023; Huang et al. 2023; Peng et al. 2024; Kim et al. 2024), which integrate unimodal expert diffusion models through dynamic weighting during the inference process. However, these approaches typically process multiple modalities in isolation or employ simplistic feature fusion strategies, resulting in suboptimal multimodal feature interaction—particularly in scenarios that demand collaborative optimization between visual masks and textual descriptions.
@@ 2.1 Diffusion Models for Image Generation
Recent advancements in generative modeling have witnessed a paradigm shift from GAN-based architectures (You et al. 2022; Du et al. 2023; Che Azemin et al. 2024) to diffusion models (Po et al. 2024; He et al. 2025; Chang et al. 2025; Huang et al. 2025), driven by their superior training stability and output diversity. Early diffusion models like DDPM (Ho et al. 2020) and DDIM (Song et al. 2020) established foundational sampling pipelines, while subsequent works focused on computational efficiency (Rombach et al. 2022; Xue et al. 2024; Chadebec et al. 2025) and conditional control (Podell et al. 2023; Yoon et al. 2023; Peng et al. 2024). Latent diffusion models (LDM) (Rombach et al. 2022) reduced computational costs by operating in compressed latent spaces, enabling high-resolution synthesis. ControlNet (Zhang et al. 2023), ControlNeXt (Peng et al. 2024), and T2I-Adapter (Mou et al. 2024) further enhanced controllability by introducing external conditioning networks, though their multimodal integration is still confined to the addition of independent feature maps.
The integration of vision transformers has driven recent breakthroughs, where diffusion transformers (DiTs) (Peebles and Xie 2023; Wang et al. 2025b) demonstrate superior performance over U-Net architectures (Ronneberger et al. 2015) through global attention mechanisms, with representative works including FLUX.1 (Labs 2024), OminiControl (Tan et al. 2024, 2025), and OmniGen (Xiao et al. 2025). However, these methods often incur significantly higher computational costs.
@@ 2.2 Facial Image Synthesis with Multimodal Conditioning
Facial synthesis techniques (Kim et al. 2023; Melnik et al. 2024; Song et al. 2025) have evolved from unimodal control paradigms to multimodal control paradigms, with recent research efforts focusing on the collaborative integration of various modalities, such as text and semantic masks (Du et al. 2023; Meng et al. 2025). Methods like TediGAN (Xia et al. 2021) and MM2Latent (Meng et al. 2025) leverage the disentangled properties of the $\mathcal{W}$ latent space in StyleGANs to achieve multimodal-driven facial synthesis by fusing latent vectors from multiple modalities. In contrast, diffusion-based approaches, such as GCDP (Park et al. 2023), have aimed to enhance semantic alignment in synthesized faces; however, they are limited by the restrictive assumptions of joint Gaussian distributions. Additionally, methods like UaC (Nair et al. 2023) and CoDiffusion (Huang et al. 2023) dynamically integrate multiple unimodal expert models during inference, which results in a more than two-fold increase in computational cost. Meanwhile, DDG (Kim et al. 2024) combines diffusion models with GANs, utilizing ControlNet (Zhang et al. 2023; Peng et al. 2024) for semantic mask and text multimodal feature fusion. Still, the expressiveness of the StyleGAN latent space representation constrains the quality of synthesized faces.
@@ 3.1 Preliminaries
Diffusion transformer (DiT) (Peebles and Xie 2023) successfully breaks through the limitations of local perception by adopting a pure transformer architecture to replace the traditional U-Net backbone network. This achieves an integration of global dependency modeling and multimodal information fusion, thereby significantly enhancing the model’s capability to model semantic consistency in complex scenarios. At time step $t$, the model utilizes stacked tansformer blocks to synchronously process noisy image tokens $\mathbf{X}_{t}\in\mathbb{R}^{N\times d}$ and text condition tokens $\mathbf{C}_{T}\in\mathbb{R}^{L\times d}$. Each transformer block incorporates a multi-head attention computation via Equation 1, ensuring in-depth interaction and fusion between visual and semantic information.
@@
Rotary Positional Embedding. To overcome the translation invariance limitation of traditional absolute position encoding, DiT introduces Rotary Position Encoding (RoPE) (Heo et al. 2024) to explicitly model spatial relationships. Each visual token at position $(i,j)$ is modulated via:
@@ 3.2 Model Design
Figure 1 illustrates the overall framework of the proposed MDiTFace method, which aims to synthesize high-fidelity facial images under the collaborative constraints of textual descriptions and semantic masks. Built upon the FLUX.1 text-to-image generation model (Labs 2024; Labs et al. 2025), MDiTFace employs a unified tokenization strategy to process heterogeneous mask-text conditional inputs, projecting them through modality-specific embedders into a shared latent space where the noise image tokens $\mathbf{X}_{t}$ resides, thereby eliminating representation disparities across heterogeneous modalities. To achieve synchronous processing and achieve a thorough fusion of multiple conditional token sequences, which include mask tokens $C_{M}$ and text tokens $C_{T}$, we have conducted a comprehensive redesign and subsequent implementation of the multivariate transformer block. These blocks incorporate an innovatively designed decoupled attention mechanism that enables flexible interaction among multimodal tokens while partitioning internal computations into dynamic and static pathways. By eliminating the implicit dependency between mask tokens $C_{M}$ and temporal embeddings, the relevant features of the static pathway can be cached after the initial computation and reused across denoising steps, thereby significantly reducing redundant computational overhead.
@@
However, experimental results indicate that this simple fusion strategy leads to poor mask-condition consistency in synthesized facial images (see Table 3 for details). To address this issue, our proposed MDiTFace method employs a unified tokenization strategy to process multimodal input conditions: The mask is encoded by reusing the VAE encoder, preserving spatial position information while maintaining architectural simplicity. The text is encoded using the pre-trained T5 encoder (Raffel et al. 2020) to capture semantic richness. Subsequently, the encoded features from both modalities are projected into a latent space shared with the noisy image tokens $\mathbf{X}_{t}$ via modality-specific embedders, bridging representation gaps between heterogeneous modalities. This transformation is formalized in Equation 4:
@@ 3.4 Training Strategy
Training Objective. We employ a lightweight adaptation architecture, which only requires low-rank adaptation (LoRA) (Hu et al. 2021) fine-tuning on the visual embedder and some linear layers within the multivariate transformer blocks. This approach results in training fewer than 0.1% additional parameters, thereby significantly reducing the computational cost of training. The model is optimized using the flow matching loss, with the specific formula as follows:
@@ 4 Experiments
Comparisons TediGAN (2021) GAN 0.7130 0.4885 1.562 87.89 23.90 2.46 -0.1446
UaC(2023) Diffusion 0.6027 0.5761 1.982 84.71 25.52 3.41 -0.4001
PixelFace+ (2023) GAN 0.5720 0.5640 1.273 93.82 26.16 2.61 0.4608
CoDiffusion (2023) Diffusion 0.7381 0.5845 0.734 83.75 24.51 3.22 -0.1265
ControlNeXt (2024) Diffusion 0.8037 0.5469 1.042 91.65 25.88 2.66 0.3675
MM2Latent (2025) GAN 0.8252 0.5124 1.178 85.05 24.61 3.05 0.2754
@@ 4.1 Experimental Setup
Datasets. MM-CelebA (Lee et al. 2020) is a large-scale, multimodal facial synthesis benchmark dataset widely used for evaluation. It comprises 30,000 highly diverse, high-resolution real-world facial images. Each image is accompanied by 10 distinct textual descriptions and a manually annotated semantic mask. The semantic masks cover 19 semantic categories, including major facial components such as hair, skin, eyes, and nose, as well as accessory categories like glasses and clothing. For model training and testing, the MM-CelebA dataset is split into training and test sets in a 9:1 ratio. In addition, to further evaluate the generalization capability of our model, we extended the FFHQ-Text dataset (Zhou 2021) into a mask-text multimodal version, denoted as MM-FFHQ, where the semantic masks were obtained using a pre-trained facial parser (Zheng et al. 2022).
Metrics. We employ a range of widely recognized metrics to evaluate the performance of our model, spanning image quality, conditional consistency, and human preference. For image quality,we employ the no-reference TOPIQ metric (Chen et al. 2024a) to evaluate the overall quality of synthesized facial images. Additionally, we utilize the CLIP Maximum Mean Discrepancy (CMMD) to quantify the distributional divergence between synthesized and real facial images in the CLIP feature space (Jayasumana et al. 2024), and the Learned Perceptual Image Patch Similarity (LPIPS) to assess perceptual distortion between synthesized and original images (Zhang et al. 2018). For conditional consistency, we measure semantic alignment between text and images using the CLIP score (Radford et al. 2021), evaluate pixel-level alignment between segmentation masks and images via Mask Accuracy, and quantify structural similarity between synthesized and real facial images using the DINO Structure Distance (DSD) (Tumanyan et al. 2022). Additionally, we employ a pre-trained aesthetic scoring model (Xu et al. 2023) to compute the Image Reward Score (IRS), which serves as an indirect proxy for the human-centric visual appeal of the generated images.
Implementation Details. All experiments were conducted on a server equipped with 8 NVIDIA A100 GPUs, each with 80 GB of memory. We employed parameter-efficient LoRA fine-tuning with a rank of 8, and selected Prodigy (Mishchenko and Defazio 2024) as the optimizer. The base learning rate was set to 1.0, and the weight decay coefficient was configured as 0.01. During training, we used a batch size of 16 and optimized the model for 5,000 steps. For inference, we employed the Flow-Matching Euler Discrete scheduler, with a random seed of 42 and 28 sampling steps.
@@ 4.2 Quantitative Evaluation
Comparisons TediGAN (2021) GAN 0.6992 0.6420 1.091 86.81 25.03 3.32 -0.4847
UaC (2023) Diffusion 0.6106 0.6752 1.796 86.17 26.97 6.38 -0.8851
PixelFace+ (2023) GAN 0.5609 0.6705 1.917 95.02 26.60 3.08 -0.2616
CoDiffusion (2023) Diffusion 0.5035 0.6767 2.178 81.78 23.03 3.85 -1.2101
ControlNeXt (2024) Diffusion 0.7594 0.6165 1.073 86.52 27.80 3.84 -0.4063
MM2Latent (2025) GAN 0.8198 0.6371 0.437 84.24 26.37 3.52 -0.1197
@@ 4.3 Qualitative Comparison
Figure 3 displays the facial images synthesized by the proposed MDiTFace and competing methods under the same multimodal conditions. It is evident that the facial images synthesized by MDiTFace exhibit higher fidelity. Meanwhile, in terms of fine-grained attributes such as earrings, hair color, and hat shape, MDiTFace can better align with the input conditions, as shown in rows 1-3. This is consistent with the results of quantitative evaluation experiments. In contrast, existing methods suffer from various deficiencies. For instance, methods like UaC (Nair et al. 2023), PixelFace+ (Du et al. 2023), and CoDiffusion (Huang et al. 2023) generate facial images with noticeably inadequate fidelity, resulting in unsatisfactory image quality. Other methods, such as TediGAN (Xia et al. 2021), ControlNeXt (Peng et al. 2024), and MM2Latent (Meng et al. 2025), encounter challenging issues of attribute loss or drift, leading to discrepancies between the synthesized facial images and the expected input conditions. Additionally, when handling the task of facial synthesis under extreme poses, as depicted in row 4, MDiTFace still maintains significant robustness compared to existing methods. More examples can be found in the Supplementary Material.